OpenStudy (anonymous):

If x 3^ = -27, then what is x? -9 -3 3 9

OpenStudy (anonymous):

-3 c:

OpenStudy (solomonzelman):

x^3 = -27 x^3=(-3)^3 x=-3

OpenStudy (anonymous):

-3 is the answer

OpenStudy (anonymous):

Thanks guys. Try this. Match the following items by evaluating the expression for x = -2. x -2^ x -1^ x 0^ x 1^ x 2^